OpenStudy (anonymous):

A plumber charges $50 to come to a house and $20 an hour for work time. If h represents the number of hours worked, which expression gives the total amount the plumber charges?

OpenStudy (imstuck):

You know right off that no matter how much time the plumber spends working on your toilet, he is going to charge $50 for the visit. That $50 is non-negotiable. 1 hours' labor or 100 hours' labor, he is making $50 for walking in your door. What makes him his money is the number of hours he actually spends working on your toilet, because he charges $20 per hour. So the number of hours he works on your toilet times $20 per hour is added to his $50 non-negotiable fee to give us an expression of 50 + 20h

OpenStudy (imstuck):

or 20h + 50 works too, of course.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your asome

OpenStudy (anonymous):

can you help anser this Which phrase could be represented by this expression? (z-7) ----- 4

OpenStudy (imstuck):

what are the phrases, assuming they gave you some.

OpenStudy (anonymous):

A. seven less than one-fourth of a number B. four times the difference of a number and seven C. one-fourth the difference of a number and seven D. seven less than four times a number

OpenStudy (imstuck):

Without anything to go on, I would say "the quantity 7 less than z divided by 4"

OpenStudy (anonymous):

those are the opens

OpenStudy (imstuck):

c

OpenStudy (anonymous):

thx you are the best helper

OpenStudy (anonymous):

Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

as you can tell im bad at math

OpenStudy (imstuck):

working on it,,,just give me a sec.

OpenStudy (anonymous):

and my mom is a sleep so she cant help that is why im do this

OpenStudy (anonymous):

and how do you change your piter

OpenStudy (imstuck):

That's ok...no need for an explanation! We will call Sharon's age x. If Anne is 3 years older than Sharon, then Anne's age is x + 3. If Tom is 5 years older than Anne, his age is 5 + (x + 3). They all add up to equal 62. So let's do that...add them all together and set them equal to 62.

OpenStudy (imstuck):

x + (x + 3) + [5 + (x + 3)] = 62 which simplifies beautifully down to x + x + 3 + 5 + x + 3 = 62 and even further to 3x + 11 = 62. Solving for 3x we get 3x = 62 - 11 or 3x = 51 Solving for x we get x = 17. Sharon is 17, Anne is 20, and Tom is 25

OpenStudy (imstuck):

You can change your picture by going up to the upper right hand of the page and clicking "Settings". That should allow you to change your profile pic.